You need to make a buffer solution with pH of 8.0. You have the following reagents on your shelf:

HC3​H5O3​    Ka=1.4*10-4 NaF
HF    Ka=6.6*10-4 Tris-base
HC2H3O2   ​    Ka=1.8*10​-5 NH​3
HClO    Ka=2.9*10​-8 NaC2H3O2
Tris-HCl Ka=8.3*10​-9 NaClO
NH4​Cl    Ka=5.6*10-10 NaC3H5O3

a) Which reagents would you use to make the buffer solution. Briefly​ explain why you would choose these.

b) In what ration of molar concentrations would you combine them?
Answer #1

a) According to Handerson Hasselbach equation for buffers,.

pH = pKa + log [base] / [acid]

And, buffer would be more effective when pH = pKa

ie when [base] = [acid]

So,

a)

Tris HCl which is having

pKa = - log (Ka) = - log (8.3 x 10^-9) = 8.0 (approx)

and its corresponding base ie Tris base would be best suitable to make the buffer of desired pH.

other reagents would have pKa values that have a higher mismatch with that of pH value ie 8.

b)

since, pH = pKa which means [base] = [acid] . Therefore we would combine them in 1:1 mole ratio.